OpenStudy (anonymous):

Simplify the expression. (-2 - 2i)(-4 + 6i) @nikato

OpenStudy (anonymous):

does "simplify" mean "multiply"?

OpenStudy (anonymous):

8 - 4i 8 - 12i 20 - 4i -4 - 4i Im not sure :/

OpenStudy (anonymous):

\[(a+bi)(c+di)=(ac-bd)+(ad+bc)i\]

OpenStudy (anonymous):

or just multiply out like in algebra "foil" (ugh) and when you get \(i^2\) replace it by \(-1\)

OpenStudy (anonymous):

So it would be the fourth option sinc it contains a negative @satellite73 ?

OpenStudy (triciaal):

@satellite73 no it doesn't to simplify means to reduce to its lowest form in this problem means to get rid of parenthesis and to do that we need to multiply

OpenStudy (anonymous):

Okay @triciaal so what do I do?

OpenStudy (triciaal):

for this problem yes you need to multiply the point I am making is that sometimes simplify may mean you have to divide etc whatever is necessary to reduce to the lowest form

OpenStudy (anonymous):

Okay.. so steps are?

OpenStudy (triciaal):

what @satellite73 has above is correct

OpenStudy (anonymous):

Could you help with that process?

OpenStudy (anonymous):

what i was trying to say was that the instructions are incorrect there is no such mathematical operation as "simplify" in this case it means "multiply"

jimthompson5910 (jim_thompson5910):

Instead of memorizing a formula, you can use a table like this to expand out the expression  keep in mind that i^2 = -1